# Av. Diagonal Las Torres, 2640 Peñalolen, Chile

### v1

@@ 1.1 EE in odd dimensions, shape dependence and holography
allows one to compute the EE as the area of the RT surface $\mathbf{A}\left(\Sigma_{A}\right)$, a minimal surface in the bulk which is homologous to the entangling region defined in the anti de Sitter (AdS) boundary, where the CFT is defined. Using this, one can obtain expansions of the form (1) for arbitrary boundary regions, where $\delta$ is a geometric regulator along the holographic direction. With some more work, and in the particular case of $d=3$, it is possible to write an explicit geometric formula for the universal term $F(A)$. In the vacuum state, this reads Alexakis:2010zz ; Fonda:2015nma ; Anastasiou:2020smm
@@
In this case, the doubled RT surface $2\Sigma_{A}$ is embedded in $\mathbb{R}^{5}$. Just like in the three-dimensional case, this expression is, by construction, free of UV divergences and can be used to evaluate $F(A)$ for $d=5$ holographic Einstein gravity for general regions once the corresponding RT surface has been determined. From this expression, which is novel in the physics literature, [^2] On the other hand, $\mathbf{W}_{5}$ has previously been derived using different methods in the mathematical literature. In that context, the construction of $\mathbf{W}_{5}$ has included various papers with conflicting results which have finally converged to a functional which agrees with the one presented in Eq. (6) — see Refs. guven2005conformally ; zhang2017grahamwittens ; Graham:2017bew ; Gover:2016buc ; Olanipekun:2021htq ; martino2023duality and references therein. we can derive a number of general results. On the one hand, it follows that Eq.  (4) does not go through to the five-dimensional case. Namely, at least for holographic theories, $F(A)$ does not have a sign and, as we show explicitly below, it can take positive, negative and vanishing values for different entangling regions. In particular, while the round ball $\mathbb{B}^{4}$ remains a local minimum — which holds true for general theories due to Mezei’s formula Mezei:2014zla ; Faulkner:2015csl — it is possible to find continuous families of regions for which $F(A)$ first grows as one deforms the ball, then it reaches a maximum, then it vanishes again for some other region and then it takes arbitrarily negative values — see Figure 5. As we explain below, a close look at the five-dimensional free-scalar and free-fermion results for $F(A)$ available in the literature reveals that $F(A)$ does not have a sign in those cases either, so it is reasonable to expect this to be a general feature of $d=5$ CFTs.
Our derivation of the above formula for $\mathbf{W}_{5}$ departs from the techniques utilized in Refs. Alexakis:2010zz ; Fonda:2015nma ; Anastasiou:2020smm for the derivation of $\mathbf{W}_{3}$ in the holographic context. It relies on the so-called Conformal Renormalization method, which we explain in some detail in the following subsection.
@@ 1.2 Conformal Renormalization and holographic EE
Because of the geometric properties of asymptotically AdS (AAdS) manifolds, the gravitational on-shell action as well as any other local functional of boundary-anchored hypersurfaces — such as the holograhpic EE — are divergent. As such, in order to define finite functionals, one needs a renormalization prescription. In Refs. Henningson:1998gx ; deHaro:2000vlm ; Bianchi:2001kw ; Skenderis:2002wp , the prescription of holographic renormalization (HR) was developed, wherein a series of boundary counterterms which are defined in terms of the induced metric and its Riemannian curvature are added at the AdS boundary, such that the on-shell gravitational action is rendered finite and its variational principle is made into a well-defined Dirichlet boundary problem. Later, in Refs. Maldacena:2011mk ; Anastasiou:2016jix , it was shown that the HR counterterms for Einstein-AdS gravity can be obtained asymptotically at the AdS boundary by embedding the theory into Conformal Gravity (CG). This is possible in four bulk dimensions as every solution of Einstein gravity with or without cosmological constant — is also a solution of CG, and also, for Einstein-AdS spacetimes, the CG on-shell action is equal to the renormalized on-shell action of Einstein-AdS gravity, written in Macdowell-Mansouri form Miskovic:2009bm . Also, in Ref. Grumiller:2013mxa it was shown that the Weyl-squared action of CG is finite when evaluated for any four-dimensional AAdS manifold.
@@
In the case of six-dimensional AAdS manifolds, Lü, Pang and Pope (LPP) have shown that there is a unique combination of the three point-wise conformal invariants in six dimensions which admits the Schwarzschild-AdS black hole as a solution Lu:2011ks ; Lu:2013hx . Interestingly, all Einstein spaces are solutions of this same linear combination of conformal invariants Anastasiou:2020mik ; Anastasiou:2023oro . We shall refer to this six-dimensional version of CG as “LPP CG”. Furthermore, when the LPP CG action is evaluated in Einstein-AdS spacetimes, then renders finite since it reduce to the renormalized Einstein-AdS action Anastasiou:2020mik . This procedure for renormalizing Einstein-AdS gravity by embedding it into CG was dubbed Conformal Renormalization.
@@
In the computation of holographic EE, the Conformal Renormalization prescription provides a natural way to isolate the finite term in odd-dimensional CFTs dual to Einstein-AdS. This is because the EE can be computed directly from the gravitational on-shell action, using the generalized gravitational entropy formula Lewkowycz:2013nqa . The finiteness of the latter gets inherited by the former. Following this idea, in Ref. Anastasiou:2022ljq the holographic EE functional for Einstein-AdS gravity in four bulk dimensions was derived starting from CG. This is achieved using the generalized gravitational entropy formula on the CG action, which isevaluated on the conically-singular orbifold obtained through the replica trick Calabrese:2004eu and using the relations given in Ref. Fursaev:2013fta . Then, the resulting functional was identified with the integrand of the Graham-Witten anomaly Graham:1999pm , which corresponds to a pointwise conformally invariant functional defined on the codimension-two hypersurface localized at the conical singularity. This functional was explicitly used to derive not only the renormalized holographic EE of Einstein-AdS, but also the so-called “reduced Hawking mass” and Willmore energy functionals which, in other contexts, are related to interesting quantities such as the entanglement susceptibility and to global bounds on information Fonda:2015nma .
@@
The remainder of the paper is as follows. In Sec. 2 we review the derivation of Willmore energy $\mathbf{W}_{3}$ and reduced Hawking mass $\mathbf{I}_{3}$ from evaluating the four-dimensional CG action in a manifold with a conical defect. This recipe generates a codimension-two functional $\mathbf{L}$, known as Graham-Witten anomaly, from which $\mathbf{W}_{3}$ and $\mathbf{I}_{3}$ appear as particular cases. In Sec. 3 we extend the derivation presented in the previous section to the six-dimensional LPP CG action, which, after evaluating in the conically singular manifold, produces a four-dimensional functional $\mathbf{F}$, which coincides with the Graham-Reichert anomaly. We observe that the analogous particular cases to the ones considered in two dimensions less allow us to identify a generalized Willmore energy $\mathbf{W}_{5}$ which matches proposals in the mathematical literature as well as a “generalized reduced Hawking mass” $\mathbf{I}_{5}$. We show that the relation between holographic EE in $d=3$ and Willmore energy $\mathbf{W}_{3}$ by means of the doubling of the RT surface also holds for holographic EE in $d=5$ and the novel $\mathbf{W}_{5}$ functional. We perform explicit checks of this relation for entangling regions consisting of round balls, slightly deformed balls and strips, obtaining the same results from both expressions. In Sec. 4 we study the global shape-dependence of the $\mathbf{W}_{5}$ functional and observe that this quantity is neither bounded from below nor from above. Related observations in the case of free fields lead us to conjecture that $F(A)$ is unbounded both from below and from above for general five-dimensional CFTs. In Sec. 5 we conclude with some general comments and future directions in light of the results obtained in this paper. Our notation and conventions are summarized in Appendix A. Certain intermediate calculations explained in the bulk of the paper appear in Appendices C, D and B.
@@ 2 Holographic EE in $d=3$ as Willmore energy
The link between holographic EE and Willmore energy can be understood from the prescription described in the following subsections. First, consider the RT surface ${\Sigma_{A}}$ associated to the entangling region $A$, this is, possessing the same boundary $\partial A=\partial{\Sigma_{A}}$, homologous and with minimal area. Now, glue an identical copy ${\Sigma_{A}}^{\prime}$ along its boundary, i.e. $\partial{\Sigma_{A}}=\partial{\Sigma_{A}}^{\prime}$. The obtained doubled-copied submanifold $2{\Sigma_{A}}={\Sigma_{A}}\cup{\Sigma_{A}}^{\prime}$ is a closed surface embeddable (through a Weyl rescalling of the ambient space) in $\mathbb{R}^{3}$ to which we can associate a Willmore energy [^4] This relation only works for entangling regions in the ground state of the CFT, which is dual to pure AdS. As the Poincaré patch of a constant-time slice of AdS4 is conformally equivalent to $\mathbb{R}^{3}/\mathbb{Z}_{2}$, the procedure considers doubling the bulk across the boundary, and then performing a conformal transformation to obtain $\mathbb{R}^{3}$. The RT surface is therefore also doubled. — in Figure 2 below the doubling of the RT surface is presented in the case of a spherical entangling surface. Based on this, the finite piece of EE can be expressed as Babich:1992mc ; Alexakis:2010zz ; Fonda:2015nma ; Anastasiou:2020smm
@@ 2.1 Energy functionals from CG in four dimensions
$\int_{\mathcal{M}}\text{d}^{4}x\sqrt{|g|}\,\mathcal{X}_{4}=32\pi^{2}\chi\left(\mathcal{M}\right)+\int_{\partial\mathcal{M}}\text{d}^{3}x\sqrt{|h|}\,\mathcal{B}_{3}\,,$ (15)
@@
This notion of curvature suggests a link to a gravity theory where the (full) Weyl tensor plays an essential role, that is, CG. At the fundamental level, the mechanism of embedding an Einstein action supplemented with a topological term in a CG, denoted conformal covariantization, is far from rigorously defined. However, we will see that it allows to derive the correct codimension-two energy functional. In this line, in order to c.c. $\mathcal{Y}_{4}\big{|}_{\text{E}}$ we just complete it to full CG, this is
@@ 2.2 Renormalized area
Since the area term in the RT formula (2) comes from the conical contribution of the Einstein-Hilbert action due to Eq. (12), then, at the saddle point, the surfaces minimizing the area should be embedded into Einstein spacetimes. As a consequence, it is expected that renormalized area should result from the same class of ambient spacetimes but with AAdS asymptotics. Even though, the parent action of $\mathcal{K}_{\Sigma}$ is a higher-curvature gravity theory,i.e, Conformal Gravity, this admits an Einstein sector in its set of solutions Anastasiou:2020mik ; Anastasiou:2023oro . For them, the Weyl tensor acquires the particularly simple expression of Eq. (17) for which we can exchange the Riemann tensor of the ambient space with quantities defined on the embedded surface using the Gauss-Codazzi relation
@@ 2.4 Willmore energy
in which, in the case of a constant-time slice of Euclidean Poincaré-AdS space, $g_{\alpha\beta}\text{d}x^{\alpha}\text{d}x^{\beta}=\frac{L_{\star}^{2}}{z^{2}}\left(\text{d}z^{2}+\text{d}\mathbf{x}^{2}\right)$, with $\mathbf{x}=(x_{1},x_{2})$, we choose $\varphi=-\log\frac{z}{L_{\star}}$ to remove the conformal factor, arriving to the three-dimensional Euclidean space
@@ 3.1 CG in six dimensions
This combination renders the Einstein-Hilbert action finite for solutions whose boundary is conformally flat Anastasiou:2020zwc . The tilde in the above functional makes reference to the fact that the finiteness is not achieved for an arbitrary AAdS geometry, but limited to the class just mentioned. As a consequence, additional counterterms on top of the topological term $\mathcal{X}_{6}$ are required. At this point, it is difficult to think of an underlying principle which could give rise to such correction while also reproducing the topological term.
The proposal of Conformal Renormalization considers a symmetry enhancement at the level of the action: from general diffeomorphism invariance to Weyl invariance Anastasiou:2020mik . This feature becomes manifest by the vanishing of the local part of the action in Eq. (42) for global AdS spaces. As a consequence, the action is factorizable by the AdS curvature or, in other words, a given polynomial of the Weyl tensor for Einstein spaces — a quantity that vanishes for AdS constant curvature configurations —, Miskovic:2014zja ; Anastasiou:2018mfk
@@
On the other hand, the quadratic combination $\mathcal{Y}_{4}\big{|}_{\text{E}}$ in the polynomial cannot be directly related to conformal invariants in six dimensions. By itself, it can be cast as a six-derivative object by introducing the Schouten tensor. However, to restore Weyl covariance, we need to supplement it with a Cotton-squared $C^{2}=C_{\alpha\beta\gamma}C^{\alpha\beta\gamma}$ (which vanishes for Einstein metrics) and a surface term $\hat{J}^{\alpha}\equiv 8W^{\alpha\gamma\delta\beta}C_{\gamma\lambda\beta}-W^{\gamma\delta}_{\beta\varepsilon}\nabla^{\alpha}W^{\beta\varepsilon}_{\gamma\delta}$. Taking all this into account, we have [^8] For more details on the conformal covariantization of action (43), we refer the reader to Ref. Anastasiou:2020mik .
@@
where $J_{\alpha}\big{|}_{\text{E}}\equiv\frac{1}{2}\nabla_{\alpha}\left(W^{\beta\gamma}_{\delta\kappa}\big{|}_{\text{E}}W^{\delta\kappa}_{\beta\gamma}\big{|}_{\text{E}}\right)$. The connection between CG — in the particular form of LPP action — and the Einstein sector of the theory is made manifest by a suitable choice of the coupling, $\alpha=-L_{\star}^{4}/(384\pi G_{\text{N}})$. Interestingly, the resulting boundary term renders the action of Eq. (55) finite for Einstein-AdS spacetimes with a generic boundary geometry, namely, the renormalized Einstein action Anastasiou:2020mik .

### v2

@@ 1.1 EE in odd dimensions, shape dependence and holography
allows one to compute the EE as the area of the RT surface $\mathbf{A}\left(\Sigma_{A}\right)$, a minimal surface in the bulk which is homologous to the entangling region defined in the anti de Sitter (AdS) boundary, where the CFT is defined. Using this, one can obtain expansions of the form (1) for arbitrary boundary regions, where $\delta$ is a geometric regulator along the holographic direction. With some more work, and in the particular case of $d=3$, it is possible to write an explicit geometric formula for the universal term $F(A)$. In the vacuum state, this reads Babich:1992mc ; Alexakis:2010zz ; Fonda:2015nma ; Anastasiou:2020smm
@@
In this case, the doubled RT surface $2\Sigma_{A}$ is embedded in $\mathbb{R}^{5}$. Just like in the three-dimensional case, this expression is, by construction, free of UV divergences and can be used to evaluate $F(A)$ for $d=5$ holographic Einstein gravity for general regions once the corresponding RT surface has been determined. From this expression, which is novel in the physics literature, [^2] On the other hand, $\mathbf{W}_{5}$ has previously been derived using different methods in the mathematical literature. In that context, the construction of $\mathbf{W}_{5}$ has included various papers with conflicting results which have finally converged to a functional which agrees with the one presented in Eq. (6) — see Refs. guven2005conformally ; zhang2017grahamwittens ; Graham:2017bew ; Gover:2016buc ; Blitz:2021qbp ; Olanipekun:2021htq ; martino2023duality and references therein. we can derive a number of general results. On the one hand, it follows that Eq.  (4) does not go through to the five-dimensional case. Namely, at least for holographic theories, $F(A)$ does not have a sign and, as we show explicitly below, it can take positive, negative and vanishing values for different entangling regions. In particular, while the round ball $\mathbb{B}^{4}$ remains a local minimum — which holds true for general theories due to Mezei’s formula Mezei:2014zla ; Faulkner:2015csl — it is possible to find continuous families of regions for which $F(A)$ first grows as one deforms the ball, then it reaches a maximum, then it vanishes again for some other region and then it takes arbitrarily negative values — see Figure 5. As we explain below, a close look at the five-dimensional free-scalar and free-fermion results for $F(A)$ available in the literature reveals that $F(A)$ does not have a sign in those cases either, so it is reasonable to expect this to be a general feature of $d=5$ CFTs.
Our derivation of the above formula for $\mathbf{W}_{5}$ departs from the techniques utilized in Refs. Babich:1992mc ; Alexakis:2010zz ; Fonda:2015nma ; Anastasiou:2020smm for the derivation of $\mathbf{W}_{3}$ in the holographic context. It relies on the so-called Conformal Renormalization method, which we explain in some detail in the following subsection.
@@ 1.2 Conformal Renormalization and holographic EE
Because of the geometric properties of asymptotically AdS (AAdS) manifolds, the gravitational on-shell action as well as any other local functional of boundary-anchored hypersurfaces — such as the holograhpic EE — are divergent. As such, in order to define finite functionals, one needs a renormalization prescription. In Refs. Henningson:1998gx ; deHaro:2000vlm ; Bianchi:2001kw ; Skenderis:2002wp , the prescription of holographic renormalization (HR) was developed, wherein a series of boundary counterterms which are defined in terms of the induced metric and its Riemannian curvature are added at the AdS boundary, such that the on-shell gravitational action is rendered finite and its variational principle is made into a well-defined Dirichlet boundary problem. Later, in Refs. Maldacena:2011mk ; Anastasiou:2016jix , it was shown that the HR counterterms for Einstein-AdS gravity can be obtained asymptotically at the AdS boundary by embedding the theory into Conformal Gravity (CG). This is possible in four bulk dimensions because every solution of Einstein gravity, with or without a cosmological constant, is also a solution of CG. Moreover, for Einstein-AdS spacetimes, the CG on-shell action is equal to the renormalized on-shell action of Einstein-AdS gravity, expressed in Macdowell-Mansouri form. Miskovic:2009bm . Also, in Ref. Grumiller:2013mxa it was shown that the Weyl-squared action of CG is finite when evaluated for any four-dimensional AAdS manifold.
@@
In the case of six-dimensional AAdS manifolds, Lü, Pang and Pope (LPP) have shown that there is a unique combination of the three point-wise conformal invariants in six dimensions which admits the Schwarzschild-AdS black hole as a solution Lu:2011ks ; Lu:2013hx . Interestingly, all Einstein spaces are solutions of this same linear combination of conformal invariants Anastasiou:2020mik ; Anastasiou:2023oro . We shall refer to this six-dimensional version of CG as “LPP CG”. Furthermore, when the LPP CG action is evaluated in Einstein-AdS spacetimes, it becomes finite as it reduces to the renormalized Einstein-AdS action Anastasiou:2020mik . This procedure for renormalizing Einstein-AdS gravity by embedding it into CG was dubbed Conformal Renormalization.
@@
In the computation of holographic EE, the Conformal Renormalization prescription provides a natural way to isolate the finite term in odd-dimensional CFTs dual to Einstein-AdS. This is because the EE can be computed directly from the gravitational on-shell action, using the generalized gravitational entropy formula Lewkowycz:2013nqa . The finiteness of the latter gets inherited by the former. Following this idea, in Ref. Anastasiou:2022ljq the holographic EE functional for Einstein-AdS gravity in four bulk dimensions was derived starting from CG. This is achieved by applying the generalized gravitational entropy formula to the CG action, which is evaluated on the conically singular orbifold obtained via the replica trick Calabrese:2004eu and using the relations given in Ref. Fursaev:2013fta . Then, the resulting functional was identified with the integrand of the Graham-Witten anomaly Graham:1999pm , which corresponds to a pointwise conformally invariant functional defined on the codimension-two hypersurface localized at the conical singularity. This functional was explicitly used to derive not only the renormalized holographic EE of Einstein-AdS, but also the so-called “reduced Hawking mass” and Willmore energy functionals which, in other contexts, are related to interesting quantities such as the entanglement susceptibility and to global bounds on information Fonda:2015nma .
@@
The remainder of the paper is as follows. In Sec. 2 we review the derivation of Willmore energy $\mathbf{W}_{3}$ and reduced Hawking mass $\mathbf{I}_{3}$ from evaluating the four-dimensional CG action in a manifold with a conical defect. This recipe generates a codimension-two functional $\mathbf{L}$, known as Graham-Witten anomaly, from which $\mathbf{W}_{3}$ and $\mathbf{I}_{3}$ appear as particular cases. In Sec. 3 we extend the derivation presented in the previous section to the six-dimensional LPP CG action, which, after evaluating in the conically singular manifold, produces a four-dimensional functional $\mathbf{F}$, which coincides with the Graham-Reichert anomaly. We observe that the analogous particular cases to the ones considered in two dimensions less allow us to identify a generalized Willmore energy $\mathbf{W}_{5}$ which matches proposals in the mathematical literature as well as a “generalized reduced Hawking mass” $\mathbf{I}_{5}$. We show that the relation between holographic EE in $d=3$ and Willmore energy $\mathbf{W}_{3}$ by means of the doubling of the RT surface also holds for holographic EE in $d=5$ and the novel $\mathbf{W}_{5}$ functional. We perform explicit checks of this relation for entangling regions consisting of round balls, slightly deformed balls and strips, obtaining the same results from both expressions. In Sec. 4 we study the global shape-dependence of the $\mathbf{W}_{5}$ functional and observe that this quantity is neither bounded from below nor from above. Related observations in the case of free fields lead us to conjecture that $F(A)$ is unbounded both from below and from above for general five-dimensional CFTs. In Sec. 5 we conclude with some general comments and future directions in light of the results obtained in this paper. Our notation and conventions are summarized in Appendix A. Certain intermediate calculations explained in the bulk of the paper appear in Appendices B, C and D.
@@ 2 Holographic EE in $d=3$ as Willmore energy
The link between holographic EE and Willmore energy can be understood from the prescription described in the following subsections. First, consider the RT surface ${\Sigma_{A}}$ associated to the entangling region $A$, this is, possessing the same boundary $\partial A=\partial{\Sigma_{A}}$, homologous and with minimal area. Now, glue an identical copy ${\Sigma_{A}}^{\prime}$ along its boundary, i.e., $\partial{\Sigma_{A}}=\partial{\Sigma_{A}}^{\prime}$. The obtained doubled-copied submanifold $2{\Sigma_{A}}={\Sigma_{A}}\cup{\Sigma_{A}}^{\prime}$ is a closed surface embeddable (through a Weyl rescalling of the ambient space) in $\mathbb{R}^{3}$ to which we can associate a Willmore energy [^4] This relation only works for entangling regions in the ground state of the CFT, which is dual to pure AdS. As the Poincaré patch of a constant-time slice of AdS4 is conformally equivalent to $\mathbb{R}^{3}/\mathbb{Z}_{2}$, the procedure considers doubling the bulk across the boundary, and then performing a conformal transformation to obtain $\mathbb{R}^{3}$. The RT surface is therefore also doubled. — in Figure 2 below the doubling of the RT surface is presented in the case of a spherical entangling surface. Based on this, the finite piece of EE can be expressed as Babich:1992mc ; Alexakis:2010zz ; Fonda:2015nma ; Anastasiou:2020smm
@@ 2.1 Energy functionals from CG in four dimensions
$\int_{\mathcal{M}}\text{d}^{4}x\sqrt{|g|}\,\mathcal{X}_{4}=32\pi^{2}\chi\left(\mathcal{M}\right)+\int_{\partial\mathcal{M}}\text{d}^{3}x\,\mathcal{B}_{3}\,,$ (15)
@@
This notion of curvature suggests a link to a gravity theory where the (full) Weyl tensor plays an essential role, that is, CG. At the fundamental level, the mechanism of embedding an Einstein action supplemented with a topological term in a CG, denoted conformal covariantization (c.c.), is far from rigorously defined. However, we will see that it allows to derive the correct codimension-two energy functional. In this line, in order to c.c. $\mathcal{Y}_{4}\big{|}_{\text{E}}$ we just complete it to full CG, this is
@@ 2.2 Renormalized area
Since the area term in the RT formula (2) comes from the conical contribution of the Einstein-Hilbert action due to Eq. (12), then, at the saddle point, the surfaces minimizing the area should be embedded into Einstein spacetimes. As a consequence, it is expected that renormalized area should result from the same class of ambient spacetimes but with AAdS asymptotics. Even though, the parent action of $\mathcal{K}_{\Sigma}$ is a higher-curvature gravity theory, i.e., CG, this admits an Einstein sector in its set of solutions Anastasiou:2020mik ; Anastasiou:2023oro . For them, the Weyl tensor acquires the particularly simple expression of Eq. (17) for which we can exchange the Riemann tensor of the ambient space with quantities defined on the embedded surface using the Gauss-Codazzi relation
@@ 2.4 Willmore energy
in which, in the case of a constant-time slice of Euclidean Poincaré-AdS space, $g_{\alpha\beta}\text{d}x^{\alpha}\text{d}x^{\beta}=\frac{L_{\star}^{2}}{z^{2}}\left(\text{d}z^{2}+\text{d}\mathbf{x}^{2}\right)$, with $\mathbf{x}=(x_{1},x_{2})$. We choose $\varphi=-\log\frac{z}{L_{\star}}$ to remove the conformal factor, arriving to the three-dimensional Euclidean space
@@ 3.1 CG in six dimensions
This combination renders the Einstein-Hilbert action finite for solutions whose boundary is conformally flat Anastasiou:2020zwc . The tilde in the above functional makes reference to the fact that the finiteness is not achieved for an arbitrary AAdS geometry, but it is limited to the class just mentioned. As a consequence, additional counterterms on top of the topological term $\mathcal{X}_{6}$ are required. At this point, it is difficult to think of an underlying principle which could give rise to such correction while also reproducing the topological term.
The proposal of Conformal Renormalization considers a symmetry enhancement at the level of the action: from general diffeomorphism invariance to Weyl invariance Anastasiou:2020mik . This feature becomes manifest by the vanishing of the local part of the action in Eq. (42) for pure AdS spaces. As a consequence, the action is factorizable by the AdS curvature or, in other words, a given polynomial of the Weyl tensor for Einstein spaces — a quantity that vanishes for AdS constant curvature configurations —, Miskovic:2014zja ; Anastasiou:2018mfk
@@
On the other hand, the quadratic combination $\mathcal{Y}_{4}\big{|}_{\text{E}}$ in the polynomial cannot be directly related to conformal invariants in six dimensions. By itself, it can be cast as a six-derivative object by introducing the Schouten tensor. However, to restore Weyl covariance, we need to supplement it with a Cotton-squared $C^{2}=C_{\alpha\beta\gamma}C^{\alpha\beta\gamma}$ (which vanishes for Einstein metrics) and a surface term $\hat{J}^{\alpha}=8W^{\alpha\gamma\delta\beta}C_{\gamma\lambda\beta}-W^{\gamma\delta}_{\beta\varepsilon}\nabla^{\alpha}W^{\beta\varepsilon}_{\gamma\delta}$. Taking all this into account, we have [^8] For more details on the conformal covariantization of action (43), we refer the reader to Ref. Anastasiou:2020mik .
@@
where $J_{\alpha}\big{|}_{\text{E}}\equiv\frac{1}{2}\nabla_{\alpha}\left(W^{\beta\gamma}_{\delta\kappa}\big{|}_{\text{E}}W^{\delta\kappa}_{\beta\gamma}\big{|}_{\text{E}}\right)$. The connection between CG — in the particular form of LPP action — and the Einstein sector of the theory is made manifest by a suitable choice of the coupling, $\alpha=-L_{\star}^{4}/(384\pi G_{\text{N}})$. Interestingly, the resulting boundary term renders the action of Eq. (55) finite for Einstein-AdS spacetimes with a generic boundary geometry, namely, it recovers the renormalized Einstein action Anastasiou:2020mik .